A computer can print 12 pages in 8 seconds. At the same rate how many pages can it print in 24 minutes
Answer by richwmiller(17219)

Computers don't print. Printers print.
24 minutes= 24*60 =1440 seconds

12*24*60/8 =2160 pages in 24 minutes